An investor wishes to realize a return of 12% from a total of two investments. If he has ten thousand pesos invested at 10%, how much additional money should be invested at 16%?
Answer by robertb(5830)

0.10*10,000 + 0.16x = 0.12*(10,000 + x)
==> 1,000 + 0.16x = 1200 + 0.12x
==> 0.04x = 200
==> x = 5,000 pesos must be invested in the 16%.